# Hybrid Internal Model: A Simple and Efficient Learner for Agile Legged Locomotion

### 18 Dec 2023 v1

@@ 1 Introduction
In recent years, learning-based locomotion control methods for quadruped robots have achieved remarkable results (Hwangbo et al., 2019; Lee et al., 2020; Peng et al., 2020; Xie et al., 2021; Kumar et al., 2021; Yu et al., 2021; Margolis et al., 2022; Rudin et al., 2022; Miki et al., 2022b; Agrawal et al., 2022; Margolis & Agrawal, 2023; Wu et al., 2022), demonstrating agility beyond traditional control methods in the wild (Lee et al., 2020; Miki et al., 2022b) or on low-cost hardware platforms (Kumar et al., 2021; Caluwaerts et al., 2023; Wu et al., 2022). Powered by numerous and diverse simulated data, these algorithms allow the quadruped robots to work in real environments directly, and perform not only simple movements but complex skills such as manipulation (Fu et al., 2022; Ji et al., 2023; Cheng et al., 2023) and parkour (Zhuang et al., 2023; Caluwaerts et al., 2023).
To achieve substantial agility, it is necessary to provide the quadruped robots with accurate and complete external states (Hartley et al., 2018). However, most onboard sensors provide partial and noisy observations about proprioception and surroundings, impeding the development of robust robot systems. As a remedy, a two-stage training strategy is usually adopted by recent learning-based methods. First, an oracle policy is trained with access to external states such as terrain frictions and elevation maps in the simulator where domain randomization is applied to enhance generalizability across various terrain conditions. Subsequently, the knowledge in this policy is transferred to another policy through mimicking where external states are disabled (Miki et al., 2022a) or inferred from other information (Kumar et al., 2021). While these methods yield promising results, their performance is constrained in two ways: (1) Despite the benefits of domain randomization for simulation-to-real (sim2real) transfer, the introduced noises restrict further optimization of the model, particularly when they directly learn to regress absolute environmental parameters. (2) The mimic learning phase leads to discrepancies between the oracle and the mimicking policy, inevitably resulting in information loss. Although this strategy yields promising results, it complicates the training process and inevitably induces information loss in the mimicking learning phase, no matter whether it is achieved by imitation learning (Miki et al., 2022a) or adaptation methods (Kumar et al., 2021). Moreover, certain solutions require additional exteroceptive sensors, such as cameras (Agarwal et al., 2023) or lidar Rudin et al. (2022), whose configurations vary across different robot platforms. Therefore, it is challenging to train and deploy a policy on universal platforms.
@@ 2 Related Work
Meanwhile, there is also a different group of methods that incorporate dynamics learning to improve legged locomotion. DayDreamer (Wu et al., 2023) followed the Model-based Reinforcement learning (MBRL) framework, which uses a learned “world model” to synthesize infinite interactions; DreamWaQ (Nahrendra et al., 2023) uses a leaned representations via VAE (Kingma & Welling, 2014) to boost the performance of legged locomotion. What these methods have in common is the use of a regression objective to learn a model or representation, which requires the neural network to fit the targets perfectly. However, due to the uncertainty brought by domain randomization, the network is willing to fit the random noise, eventually resulting in representation collapse.

### 21 Dec 2023 v2

@@ 1 Introduction
In recent years, learning-based locomotion control methods for quadruped robots have achieved remarkable results (Hwangbo et al., 2019; Peng et al., 2020; Yang et al., 2020; Xie et al., 2021; Imai et al., 2022; Yu et al., 2021; Yang et al., 2022; Margolis et al., 2022; Rudin et al., 2022; Agrawal et al., 2022; Yang et al., 2023b; Margolis & Agrawal, 2023; Yang et al., 2023b), demonstrating agility beyond traditional control methods in the wild (Lee et al., 2020; Miki et al., 2022b) or on low-cost hardware platforms (Kumar et al., 2021; Wu et al., 2022). Powered by numerous and diverse simulated data, these algorithms allow the quadruped robots to work in real environments directly, and perform not only simple movements but complex skills such as manipulation (Fu et al., 2022; Ji et al., 2023; Cheng et al., 2023) and parkour (Yang et al., 2023a; Zhuang et al., 2023; Caluwaerts et al., 2023).
To achieve substantial agility, it is necessary to provide the quadruped robots with accurate and complete external states (Hartley et al., 2018). However, most onboard sensors provide partial and noisy observations about proprioception and surroundings, impeding the development of robust robot systems. As a remedy, a two-stage training strategy is usually adopted by recent learning-based methods. Firstly, an oracle policy is trained with access to external states such as terrain frictions and elevation maps in the simulator where domain randomization is applied to enhance generalizability across various terrain conditions. Subsequently, the knowledge in this policy is transferred to another policy through mimicking where external states are disabled (Miki et al., 2022a) or inferred from other information (Kumar et al., 2021). Although methods derived from this strategy yield promising results, their performance is constrained in two ways: (1) Despite the benefits of domain randomization for simulation-to-real (sim2real) transfer, the introduced noises restrict further optimization of the model, particularly when they directly learn to regress absolute environmental parameters. (2) This two-phase training paradigm complicates the training process and inevitably induces information loss in the mimicking learning phase, no matter whether it is achieved by imitation learning (Miki et al., 2022a) or adaptation methods (Kumar et al., 2021). Moreover, certain solutions require additional exteroceptive sensors, such as cameras (Agarwal et al., 2023) or lidar Rudin et al. (2022), whose configurations vary across different robot platforms. Therefore, it is challenging to train and deploy a policy on universal platforms.
@@ 2 Related Work
Meanwhile, there is also a different group of methods that incorporate dynamics learning to improve legged locomotion. DayDreamer (Wu et al., 2023) followed the Model-based Reinforcement learning (MBRL) framework, which uses a learned “world model” to synthesize infinite interactions; DreamWaQ (Nahrendra et al., 2023) uses a leaned representations via VAE (kingma2014auto) to boost the performance of legged locomotion. What these methods have in common is the use of a regression objective to learn a model or representation, which requires the neural network to fit the targets perfectly. However, due to the uncertainty brought by domain randomization, the network is willing to fit the random noise, eventually resulting in representation collapse.